OpenStudy (anonymous):

Suppose f(x) is a quadratic polynomial satisfying f(0)=1 f(4)−f(2)=2 f(6)−f(4)=3 What is the value of f(8)?

hartnn (hartnn):

let the quadratic polynomial be \(\large f(x)=ax^2+bx+c\) from f(0)=1, you can easily find the value of 'c', which is ?

OpenStudy (anonymous):

1?

hartnn (hartnn):

correct! so, f(x) = ax^2+bx+1 now can you find, f(2), f(4), f(6) ?? the goal is to use f(4)−f(2)=2 f(6)−f(4)=3 and get 2 equations in a and b, so that they can be solved simultaneously, to get the values of a and b

OpenStudy (anonymous):

ok i'm lost :'(

hartnn (hartnn):

ok, can you find f(2) ??

OpenStudy (anonymous):

um no

hartnn (hartnn):

@ziko1995 let him do that work of getting equations

OpenStudy (anonymous):

@hartnn ok :)

hartnn (hartnn):

ok, to find f(2), just plug in x= 2 in f(x) so, what u get after u put x=2 in ax^2+bx+1 ?

OpenStudy (anonymous):

4a + 4b + 1

OpenStudy (anonymous):

@killua_vongoladecimo It's 4a+2b+1 cause we haven't \[bx ^{2}\] :)

OpenStudy (anonymous):

sorry

hartnn (hartnn):

x=2 so, bx, gives 2x, right ?? so, its f(2)=4a+2b+1 similarly find f(4)

OpenStudy (anonymous):

hmm.. in f(4) - f(2) = 2 after using ax^2 + bx + 1 it will become (16a + 4b + 1) - (4a + 2b + 1) = 2 then when simplified 12a + 2b = 2 right?

hartnn (hartnn):

correct! simplify it. now do f(6)-f(4)=3

OpenStudy (anonymous):

20a + 2b = 3

OpenStudy (anonymous):

right? what's next?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

12a+2b=2 20a+2b=3 just subtract one equation from other and b gets eliminated!

OpenStudy (anonymous):

so a= 1/8 b= 1/4 ?

hartnn (hartnn):

correct! so, we got our f(x)! \(\huge f(x)=(1/8)x^2+(1/4)x+1\) to get f(8), just plug in x=8 there :)

OpenStudy (anonymous):

oh... so f(8) = 11?

hartnn (hartnn):

yes, i get 11 too :)

OpenStudy (anonymous):

oh so that's how it is.....thank you hartnn but i still got more questions haha...

hartnn (hartnn):

welcome ^_^ sure! ask :) better you close this first and ask in new post ?